OpenStudy (tate_langdon):

Can someone check my work and show work Simplify the following expression. -21÷ (-7) -21 ÷ (-7) A. 3 B. -3 C. -5 D. 5

OpenStudy (tate_langdon):

I say the answer is A

OpenStudy (tate_langdon):

@sunnnystrong

jhonyy9 (jhonyy9):

-21+(-7) = -21-7 = ?

OpenStudy (tate_langdon):

The answer is A

jhonyy9 (jhonyy9):

or there is minus 21 divide -7 ?

jhonyy9 (jhonyy9):

so than you are right with choice A. sure

OpenStudy (tate_langdon):

Can you show the work

jhonyy9 (jhonyy9):

-21 21 --- = ---- = 3 -7 7

OpenStudy (sunnnystrong):

Oh hey soo.... \[\frac{ -21 }{ -7 }=3\] You're right :D

OpenStudy (tate_langdon):

Thanks guys

jhonyy9 (jhonyy9):

yw